# Alpaca against Vicuna: Using LLMs to Uncover Memorization of LLMs

### v2

@@ 4.2 Evaluation Data
Data Domains: To curate diverse evaluation datasets, we extract samples from base model pre-training data (Llama, OLMo, and Falcon). Given Llama’s data unavailability, we replicate it using the RedPajama dataset, while Falcon’s RefinedWeb, sourced from Common Crawl, is accessible. OLMo’s pre-training data, Dolma, is directly utilized. Across five domains for Llama and six for Dolma—Github, C4, CC, Arxiv, and Books for Llama, with The Stack substituted for Github, PeS2o for Arxiv, and Reddit added for social media within Dolma—we select 15,000 samples from Llama, 3,000 from Falcon’s RefinedWeb, and 16,000 from OLMo’s domains, ensuring uniform distribution across sequence lengths.
Sequence Lengths Selection: To measure our method’s adaptability across varying sequence lengths (200, 300, and 500), we adopt a splitting ratio informed by real-world usage patterns. Drawing from the WildChat dataset analysis (Zhao et al., 2024), we allocate 33% of each sample as the prefix and the remaining 67% as the suffix, enhancing the representation of typical usage scenarios. See Appendix D for more details.
Average Over Three Sequence Lengths (200, 300, 500)
Model Method Access Github ArXiv CC C4 Books
Mem LCSP Dis Mem LCSP Dis Mem LCSP Dis Mem LCSP Dis Mem LCSP Dis
Alpaca P-S-Base B .291 .125 - .183 .112 - .190 .104 - .204 .114 - .208 .093 -
P-S-Inst B .270 .124 - .179 .112 - .155 .104 - .143 .114 - .131 .093 -
Reverse-LM B .229 .200 .864 .133 .196 .848 .113 .186 .843 .110 .181 .834 .122 .142 .865
GCG W .300 .110 .530 .178 .101 .379 .194 .090 .374 .208 .102 .321 .199 .080 .422
Ours B .322 .102 .864 .228 .108 .848 .214 .096 .830 .203 .090 .834 .221 .079 .865
Vicuna P-S-Base B .291 .125 - .183 .112 - .190 .104 - .203 .114 - .208 .093 -
P-S-Inst B .273 .125 - .213 .112 - .205 .114 - .191 .114 - .198 .093 -
Reverse-LM B .255 .200 .864 .200 .196 .848 .173 .186 .830 .173 .181 .834 .166 .142 .865
GCG W .300 .110 .530 .178 .101 .379 .194 .090 .374 .208 .102 .321 .199 .080 .422
Ours B .325 .096 .864 .232 .104 .853 .213 .092 .838 .201 .084 .841 .223 .079 .866
Tulu P-S-Base B .291 .125 - .183 .112 - .190 .104 - .203 .114 - .208 .093 -
P-S-Inst B .274 .124 - .207 .112 - .170 .106 - .137 .114 - .172 .093 -
Reverse-LM B .245 .200 .864 .153 .196 .848 .121 .186 .830 .117 .181 .834 .135 .142 .865
GCG W .300 .110 .530 .178 .101 .379 .194 .090 .374 .208 .102 .321 .199 .080 .422
Ours B .359 .104 .857 .237 .104 .851 .221 .094 .835 .210 .086 .836 .233 .079 .865
Seq Len Tulu-7B
200 P-S-Base B .315 .126 - .188 .107 - .198 .103 - .206 .111 - .225 .090 -
P-S-Inst B .298 .125 - .216 .107 - .176 .103 - .140 .111 - .188 .090 -
300 P-S-Base B .315 .126 - .188 .107 - .198 .103 - .206 .111 - .225 .090 -
P-S-Inst B .276 .124 - .209 .112 - .174 .106 - .142 .114 - .178 .095 -
500 P-S-Base B .263 .124 - .175 .117 - .179 .102 - .196 .117 - .184 .095 -
P-S-Inst B .247 .124 - .195 .117 - .159 .102 - .128 .117 - .149 .095 -
Table 1: Comparison of our method with black-box (‘B’) and white-box (’W’) baselines across pre-training data domains. Mem denotes the memorization score (ROUGE-L), $LCS_{P}$ is input prompt and suffix overlap, and Dis is optimized vs. initial prompt distance. Results are averaged over three sequence lengths on top, and for the Tulu-7B model, we show a breakdown at the bottom. The highest performance within each domain is bolded.
@@ 4.3 Baseline Methods
We compare against three methods under two access settings: white-box and black-box.
(1) Prefix-Suffix (P-S) sequence extraction method (Carlini et al., 2022; 2021): We apply a black box attack by prompting the model with the original prefix of the pre-training sample (i.e., the first $n$ tokens) and generating the model output. We call this baseline the Prefix-Suffix (P-S) method. We evaluate both the base model and instruction-tuned versions.
(2) GCG (Zou et al., 2023): We test a prominent white-box adversarial attack method for LMs. Our application of GCG uses the original prefix as the starting point for each sample; we train for thirty epochs and apply it to the base model.
(3) Reverse LM (Pfau et al., 2023): This model reverses the token order during training, predicting optimized prefixes given specific suffixes, using a Pythia-160M model trained on the deduplicated Pile dataset (Pfau et al., 2023; Biderman et al., 2023b; Gao et al., 2020).
@@ 4.4 Evaluation Metrics
Measuring Memorization/Reconstruction: We evaluate memorization using ROUGE-L, measuring the longest common subsequence between generated and original suffixes. Our approach aligns closely with the memorization score proposed by Biderman et al. (2023a), emphasizing ordered token matches between model-generated and true continuations.
Evaluating Prompt Overlap: As our method relies on building a prompt on the whole sequence, including the ground truth (suffix), we measure the overlap between the prompt and suffix. We aim to ensure that the prompt retains less or equal overlap compared to the original prefix-suffix combination. We use ROUGE-L to measure the overlap between the prompt and the suffix, which we denote as $LCS_{P}$.
5 Experimental Results
In this section, we present our main experimental results, comparing our method to the baselines, on the base and instruction-tuned models, with different data domains and lengths. We then further analyze different components of our method, along with changing our assumption in terms of access to suffixes and data points. We provide details of hyperparameters in Appendix A, a breakdown of the results in terms of sequence lengths and improvement percentages in Appendix B, a comparison of reconstructions from different models in Appendix C, and analysis of common patterns in subsection B.1. Finally, we include samples of optimized prompts and outputs in Appendix E.
5.1 LLMs memorize more than we think!
Table 1, which shows our main results alongside our baselines. Except for P-S-Base & GCG, all other numbers are reported over the instruction-tuned model as the target, whereas the P-S-Base is evaluated on the base model (Llama– hence the same value on all rows), to provide a comparison point. If we only compare P-S-Inst and P-S-Base, we would make a misleading conclusion that instruction-tuned models uncover less training data than their base counterparts. Our method, surprisingly, uncovers more memorization than all other baselines, even on the base model, showing that instruction-tuned models can reveal more pre-training data, even more than the base if prompted appropriately. On Github, for instance, the base LLama model exhibits a reconstruction Rouge-L score of .291, while Tulu shows .274, with the P-S-Inst baseline. Our method increases this to .322. These trends extend beyond Llama-based models to OLMo Groeneveld et al. (2024) Figure 2 and Falcon Penedo et al. (2023). Moreover, the same trend exists for larger sizes (13B, 30B). See Appendix B for the Falcon model and larger sizes results.
Figure 3: Comparison of our attack performance when the prompt is optimized over only the sequence prefix (partial access) versus when we have access to the entire sequence (default assumption through the paper). Notably, the performance of attacks relying solely on prefixes closely aligns with those utilizing the entire sequence across most domains, pointing at the robustness of the optimization toward partial access to the training point.
If we expand our baselines and take a closer look at the white-box GCG optimization, we observe that it uncovers more memorization than P-S (sequence-extraction) attacks by 1% on average, but it still falls short of our method. Github exhibits the most significant increase across all domains, while other domains generally show improvement over sequence-extraction-based methods in most settings. ReverseLM performs the worst, possibly due to its usage in a transferability setting from the Pythia model. Details of hyperparameters can be found in Appendix A, a breakdown of the results and improvement percentages in Appendix B, and samples of optimized prompts and outputs in Appendix E.
Measuring Overlap Between Prompts & Suffixes. Assessing prompt-response overlap is crucial to make sure that the optimized prompt doesn’t include the pre-training data in it (i.e. make sure we are not cheating). We introduce an overlap penalty (subsection 3.1) to mitigate this. Results in Table 1 consistently show our method achieving equivalent or lower overlap ($LCS_{p}$) in terms of ROUGE-L, with the prefix-suffix baseline. For example, our approach has significantly lower overlap in domains like GitHub, ensuring a fair comparison with baseline methods and demonstrating that prompts exist that are substantially different from the original pre-training prefix and can yet result in better reconstruction.
PII Identification. We assessed our method’s reconstructions to evaluate the degree of identifiable information (PII) revealed by categorizing 9,000 pre-training samples (CC, C4, Github) using regular expressions to identify various PII elements (phone, link, email, credit cards, street address, SSN, zip code). We then applied the same procedure to generate content from optimized prompts and compared results with ground truth, retrieving an average of 10.28% of PII from pre-training samples, a significant increase of 1.43 times compared to the 4.23% achieved by the prefix-suffix attack.
5.2 GPT-4 is NOT the best attacker!
We test GPT-4 as an alternative attacker to assess its effect on performance, finding Zephyr outperforms GPT-4 consistently for sequence length 200, maintaining superiority across all domains by a margin of 0.05 as shown in Figure 4. The performance gap narrows as the sequence length increases to 300, but Zephyr remains ahead. However, at length 500, GPT-4 starts to match or surpass Zephyr’s performance, especially notable in the ArXiv domain, possibly due to the increased difficulty of summarization with longer sequences.
@@ 5.2 GPT-4 is NOT the best attacker!
Figure 4: A comparison of our method’s performance using Zephyr and GPT-4 as attacker LLMs is shown for different iteration steps during optimization. We observe a consistent trend: performance increases across varying sequence lengths as optimization iterations increase, and Zephyr uncovers more memorization than GPT-4 by a small margin. The dots are averaged across five domains and three instruction-tuning models.
5.3 What if we don’t have access to the entire training sequence?
For a more stringent setup, we assume we only have access to prefixes and not suffixes. Our method still yields comparable results across most domains and models, revealing even higher reconstruction in some cases, as shown in Figure 3. However, prompts based on entire sequences uncover significantly more memorization in domains like Github and books due to token count discrepancies. These are resolved using a whitespace tokenizer to ensure sufficient context for prefix optimization, achieving performance parity.
5.4 What if we don’t use GPT-4 for the Meta-prompt initialization?
@@
In all previous experiments, we use the meta-prompts (see Section 3.2) on GPT-4 to have it create the initial prompts. However, we investigate the influence of a less powerful open-source model on the pipeline’s overall performance. Our choice is Mixtral-8x7B instruct (Jiang et al., 2024). Using Alpaca and a sequence length of 200 as an example, we demonstrate that utilizing Mixtral can still surpass the reconstruction performance over the prefix-suffix method, achieving better performance than P-S for base and instruct model by 6.12% and 12.62%, respectively, and in contrast, being worse than GPT-4 by 4.00%.
5.5 What goes on during the optimization process?
@@
The impact of iteration count: Our approach consists of two phases: sampling and refining. The former employs rejection sampling, while the latter iterates three times on the most promising prompt, offering feedback. Visualized in Figure 4, performance evolves through each optimization stage, revealing insights into the process’s impact on performance. Though modest initially due to untargeted prompts, performance steadily enhances with each iteration, reaching peak performance by the third iteration. Increasing iterations could further enhance performance but at the expense of higher computational costs.
Measuring Edit Distance: To analyze the optimization process, we measure the gap between the initial and refined prompts using normalized Levenshtein distance, aiming for notable discrepancies to underscore its impact. Across all models, domains, and sequence lengths (as shown in Table 1), the edit distance ranges from 0.80 to 0.88, indicating substantial modifications from the initial to the optimized prompt.
@@ 6 Related Work
Data Extraction: Several studies in the literature explore data extraction techniques in LLMs, initially focusing on base models. Yu et al. (2023) proposed methods such as Top-K selection, nucleus sampling, and temperature adjustments. Nasr et al. (2023) targeted instruction-tuned models, presenting a divergence attack prompting models like ChatGPT to repeat a word despite instability issues noted indefinitely. Additionally, Zhang et al. (2023) devised a model interrogation attack, selecting lower-ranked output tokens during generation to extract sensitive user data. Furthermore, Geiping et al. (2024) introduced a system prompt repeater for extracting sensitive system prompts, demonstrating notable success in potentially compromising entire applications or secrets.
JailBreaking: Emerging red-teaming methods exploit LLMs through jailbreaking techniques, aiming to coerce harmful behaviors (Shah et al., 2023; Li et al., 2023; Huang et al., 2023; Zeng et al., 2024; Mehrotra et al., 2023; Hubinger et al., 2024). These approaches disrupt safety mechanisms, prioritizing harmful responses over data confidentiality.
7 Conclusion and Limitations
In this work, we present a novel approach for analyzing discoverable memorization of pre-training data in instruction-tuned LLMs. Our empirical findings challenge prior assumptions by demonstrating that instruction-tuned models exhibit a higher level of memorization than their base models, using prompts other than the original prefix from the pre-training data. We emphasize that although our method and experiments uncover more reconstruction of pre-training data in instruction-tuned models than base, this does not mean one model or the other is memorizing/regurgitating more data or being more or less vulnerable. It only alludes to the fact that instruction-based prompts, the way we build them, uncover more of the pre-training data in instruction-tuned models. We would also like to acknowledge that our method is mainly an auditing method which requires access to some part of the training data. We encourage future work to explore other automated strategies for building prompts for data extraction, targeting both base and instruction-tuned models, using prompts and contexts other than the original training data.

### v3

@@ 4.2 Evaluation Data
Data Domains: We construct diverse evaluation datasets by sampling from several pre-training datasets used in base models. Specifically, we use Llama (replicated from RedPajama due to data unavailability), Falcon’s RefinedWeb (from Common Crawl), and OLMo’s Dolma. Llama spans five domains (C4, CC, Arxiv, Books, and Github), while Dolma covers six domains (C4, CC, Arxiv, Books, Reddit, Stack, and PeS2o). We ensure uniformity in sequence length distribution, selecting 15,000 samples from Llama, 3,000 from Falcon’s RefinedWeb, and 16,000 from OLMo’s domains.
Sequence Lengths Selection: To evaluate adaptability across different sequence lengths (200, 300, and 500), abbreviated as "seq.," we adopt a splitting ratio inspired by real-world usage patterns. Based on analysis from the WildChat dataset (Zhao et al., 2024), we divide each sample, allocating 33% as the prefix and 67% as the suffix, reflecting typical usage scenarios (see Appendix D for further details).
@@ 4.3 Baseline Methods
We compare against three methods under two access settings: white box and black box.
(1) Prefix-Suffix (P-S) sequence extraction (Carlini et al., 2022, 2021): A black-box attack where the model is prompted with the first $n$ tokens (prefix) of a pre-training sample to generate output, applied to both base and instruction-tuned models.
(2) GCG (Zou et al., 2023): A white-box adversarial attack that starts with the original prefix and is trained for thirty epochs on the base model.
(3) Reverse LM (Pfau et al., 2023): A method that reverses token order during training, using a Pythia-160M model trained on the deduplicated Pile dataset (Pfau et al., 2023; Biderman et al., 2023b; Gao et al., 2020).
@@ 4.4 Evaluation Metrics
Measuring Memorization/Reconstruction: In our evaluation, we use ROUGE-L to measure memorization by comparing the longest common subsequence between the generated and original suffixes, closely aligning with the memorization score introduced by Biderman et al. (2023a), which emphasizes ordered token matches between model-generated continuations and the true text. To evaluate prompt overlap, particularly in our analytical solution where the prompt includes the ground truth suffix, we assess the overlap between the prompt and suffix to ensure it does not exceed the overlap in the original prefix-suffix combination. We denote this overlap as $LCS_{P}$ and use ROUGE-L to quantify it.
5 Main Results
Evaluating on Instruction-Tuned LLMs. Table 1 summarizes our main findings and compares them with baselines across different pre-training data domains. Our method reveals significantly higher levels of memorization compared to traditional prefix-suffix methods. On average, our approach achieves a 5% increase in memorization, reaching up to 12% in scenarios with a sequence length of 500. For instance, GitHub & Tulu LM achieve a reconstruction Rouge-L score of 24.7% with prefix-suffix, whereas our method improves this to 35.9%. These results hold consistently across various models, including Llama-based models, OLMo Groeneveld et al. (2024), and Falcon Penedo et al. (2023), as well as larger models like 13B and 30B. Detailed results on the Falcon model and larger sizes are provided in Appendix B.
Figure 4: Comparison of our attack performance shows that optimizing prompts over partial sequence access versus full access (default assumption through the paper) shows similar results across domains. This highlights the robustness of optimizing prompts with limited sequence information.
Evaluating on Base LLMs. Figure 2 compares Base and Instruction-tuned LLMs, GCG, and our method. Comparing P-S-Inst and P-S-Base alone would misleadingly suggest that instruction-tuned models uncover less training data. However, our method uncovers more memorization than all other baselines, including the base model, showing that instruction-tuned models can reveal more pre-training data when prompted correctly. While the white-box GCG uncovers 1% more memorization than P-S attacks, it still falls short of our method. ReverseLM performs the worst due to its transferability setting from the Pythia model. For detailed results and improvement percentages, refer to Appendix B. Hyperparameter details are in Appendix A, and optimized prompts and outputs are in Appendix B. For runtime details of the proposed method and GCG, see Appendix D.
Prompt Overlap Analysis. As shown in  Table 1, consistently, our method achieves equivalent or lower overlap ($LCS_{p}$) in terms of ROUGE-L, with the prefix-suffix baseline. For example, our approach has significantly lower overlap in domains like GitHub, ensuring a fair comparison with baseline methods.
6 Alpaca Vs Vicuna In The Wild
6.1 CASE STUDY: Extraction of Copyrighted Books/Articles
We applied our prompt optimization technique to extract copyright infringements in training data, targeting excerpts from copyrighted books and articles across various models.
Evaluation Data. We used the Books3 subset from the Redpajamas dataset to assess Llama instruction-tuned LLMs and Project Gutenberg to evaluate OLMo, as detailed in subsection 4.2. Additionally, we selected 200 samples from BookMIA and 100 from New York Times articles to evaluate GPT-4o, which has previously been shown to memorize data from these models Shi et al. (2023); Grynbaum and Mac (2023).
Results. Figure 3 and Figure 4 demonstrates that our method consistently outperforms Prefix-Suffix in OLMo & Llama based models in Book domain. For GPT-4o, Although it often refuses or avoids verbatim repetition of training data, we achieved approximately 25% overlap on average—doubling the result of simply asking or continuing the text in BookMIA & NYT.
6.2 CASE STUDY: Eliciting Unlearned Harry Potter
Eldan and Russinovich (2023) introduced an unlearning technique to remove knowledge of the Harry Potter books through multiple unlearning steps on Llama-2 Touvron et al. (2023), resulting in a model that no longer retains the targeted content. Although querying the model before and after unlearning shows it has forgotten the information, we aim to assess the model’s behavior under adversarial prompts using our approach.
Evaluation Data. We sampled 300 passages from various Harry Potter books, each with a sequence length of 300 to provide sufficient context for prompt generation.
Results. Our optimized prompts elicited highly similar completions to the original text, achieving 23.6% overlap compared to 10.2% using prefix-suffix prompts. These findings suggest the unlearning technique is vulnerable to adversarial prompts that deviate from the original training context.
6.3 CASE STUDY: LLMs Refusal
OpenAI models frequently refuse to answer certain questions, particularly those that seek harmful responses, such as inquiries about illegal activities or hate, harassment, and violence. Recently, when prompted to continue a passage from a book or article, these models declined to respond.
@@ 6.3 CASE STUDY: LLMs Refusal
Figure 5: A comparison of our method’s performance using Zephyr and GPT-4 as attacker LLMs is shown for different iteration steps during optimization. We observe that the performance increases across varying sequence lengths as optimization iterations increase.
Evaluation Data. We use the same Harry Potter book subset from unlearning to assess refusal rates, with GPT-4o as the judge. GPT-4 and GPT-4o were evaluated on the prefix-suffix and our generated prompts. We also assessed overlap and ensured the completions closely matched the ground truth from our prompts.
Results. By comparing our generated prompts with the prefix-suffix method, we found that our approach bypasses filters, yielding responses for all 300 samples, while the prefix-suffix refusal rates are 13.65% for GPT-4 and 26.19% for GPT-4o. This demonstrates the robustness of our method in adversarial generation.
6.4 CASE STUDY: Predicting Memorization For Practical Attack
We previously discussed the applicability of our approach when $d_{\text{suffix}}$ is inaccessible, as often occurs in real-world scenarios. We developed a classifier to replace the ROUGE-L function in our optimization loop to address this. We outline the preference data and then investigate the classifier’s technical details.
Preference Data. We ran several iterations using the full training sequence, collecting optimized and non-optimized prompts per sample. Each iteration produced one optimized and 23 non-optimized prompts, generating 24 samples over three iterations. We unified preference data by merging sequence lengths and victims to train a single classifier per domain, and we downsampled non-optimized classes to overcome the data imbalance problem.
Technical Details. We train a single classifier for each domain, encompassing various sequence lengths and target entities, using DeBERTa-v3-large He et al. (2021) with weighted CrossEntropy loss. The model is trained on an H100 80GB GPU for 1500 steps with a batch size of 16 and a maximum sequence length of 512. The dataset of 20,000 samples is split into 80% training, 10% validation, and 10% testing.
Results. We assess the classifier’s performance using the macro F1 score across different data domains, achieving an average F1 score of 70% in distinguishing prompts that trigger memorized responses from those that do not. While the classifier’s performance may not be optimal, we consider this a significant step toward practical attacks in future work, which could be improved by integrating prefixes with questions in the Natural Language Inference (NLI) task or utilizing Direct Preference Optimization (DPO) Rafailov et al. (2024) to better align an LLM with the distribution of optimized prompts.
7 Ablation & Analysis
In this section, we conduct ablations and analyses to identify the key components that drive our method’s improvements over baseline models.
GPT-4 is NOT the best attacker. We evaluate GPT-4 as an alternative attacker and find that Zephyr consistently outperforms GPT-4 at a sequence length of 200, maintaining a margin of 0.05 across all domains, as shown in Figure 5. While the performance gap narrows at a sequence length of 300, Zephyr still leads. At 500 tokens, however, GPT-4 begins to match or exceed Zephyr, particularly in the ArXiv domain, where summarization complexity increases with longer sequences.
Initialization without Suffix. In previous experiments, we used the full training sequence, including suffixes, to test Instruction-Tuned LLMs with an overlap penalty to prevent cheating. In real-world settings, though, only prefixes are available to construct solutions. Despite this limitation, our method performs comparably or even better in some cases, as shown in Figure 4. Since full-sequence prompts have more tokens, they show increased memorization in domains such as GitHub and books. To address this, we use a whitespace tokenizer to optimize prefixes, ensuring that performance remains competitive.
Victim as an Attacker LLM. We tested whether using the victim model as an attacker impacts performance and compared it with using distinct attacker models across different pre-training domains. In prior experiments, the same model served as both the attacker and victim, but performance consistently lagged behind using Zephyr or GPT-4 as attackers. For example, with a sequence length of 200, Tulu LM as an attacker was 7.21% less effective than Zephyr, suggesting that using different attackers and sampling strategies significantly boosts performance.
@@ 7 Ablation & Analysis
Beyond GPT-4 for meta-prompt initialization. In previous experiments, we employed GPT-4 for meta-prompt generation (see Section 3.2), but we now investigate the effect of using a less powerful open-source model on overall performance. Specifically, we utilize Mixtral-8x7B instruct (Jiang et al., 2024). In cases such as Alpaca with a sequence length of 200, Mixtral outperforms the prefix-suffix method, yielding 6.12% and 12.62% better reconstruction performance for base and instruct models, respectively, although it falls 4.00% short of GPT-4.
Training Data or Common Patterns. We test our method’s ability to generalize beyond pre-training data using the BookMIA dataset Shi et al. (2023), which contains both training data members and non-members. Our method achieved a ROUGE-L score of 23.3 on training data members but only 16.7 on non-members, suggesting that our approach may lead the model to output memorized data rather than generalized information.
@@
The impact of iteration count. Our method comprises two phases: sampling and refining. In the sampling phase, we use rejection sampling to gather data, and in the refining phase, we iterate three times on the most promising prompt, providing feedback at each step. Figure 5 illustrates performance improvements through these optimization stages. Although initial gains are modest from untargeted prompts, performance steadily improves across iterations, peaking by the third round. Further iterations could enhance performance further but would come at higher computational costs.
@@ 8 Related Work
Data Extraction: Several studies have investigated data extraction techniques in LLMs. Yu et al. (2023) proposed sampling adjustments for base models. Nasr et al. (2023) focused on instruction-tuned models, demonstrating a divergence attack causing models like ChatGPT to repeat words indefinitely. Zhang et al. (2023) developed a model interrogation attack to extract sensitive data by selecting lower-ranked output tokens. Additionally, Geiping et al. (2024) introduced a system prompt repeater to extract sensitive system prompts, potentially compromising entire applications or secrets.
JailBreaking: Emerging red-teaming methods exploit LLMs through jailbreaking techniques, aiming to coerce harmful behaviors (Shah et al., 2023; Li et al., 2023; Huang et al., 2023; Zeng et al., 2024; Mehrotra et al., 2023b; Hubinger et al., 2024). These approaches disrupt safety mechanisms, prioritizing harmful responses over data confidentiality.
9 Conclusion
In this work, we introduce a new method to analyze how instruction-tuned LLMs memorize pre-training data. Our empirical findings indicate that instruction-tuned models show higher memorization levels than their base models when using prompts that are different from the original pre-training data. However, this increased memorization in instruction-tuned models does not imply that these models regurgitate more data or are more vulnerable. Instead, it suggests that constructing instruction-based prompts reveals more pre-training data in instruction-tuned models.
Limitations
We would like to acknowledge that our method is mainly an auditing method which requires access to some part of the training data. We encourage future work to explore other automated strategies for building prompts for data extraction, targeting both base and instruction-tuned models, using prompts and contexts other than the original training data.
Ethics Statement
Enhancing the privacy-preserving capabilities of LLMs is crucial, given their increasing prominence and involvement in various aspects of life. Our new attack, designed to extract memorized data from instruction-tuned LLMs, which are widely used in real-world applications, deepens our understanding of these models’ privacy limitations. By introducing this attack, we aim to advance the comprehension of memorization behaviors in different types of LLMs, encouraging future work to develop novel defense mechanisms to mitigate associated risks.
Acknowledgements
Funding support for the project activities of Niloofar Mireshghallah and Yulia Tsvetkov has been provided by the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300), NSF DMS-2134012, IARPA HIATUS via 2022-22072200003, and ONR N00014-24-1-2207. The views expressed are those of the author and do not reflect the official policy or position of the Department of Defense or the U.S. Government. For Aly M. Kassem, this research is supported by the Vector Scholarship in Artificial Intelligence, provided through the Vector Institute and Natural Sciences and Engineering Research Council of Canada (NSERC) by NSERC Discovery Grant. This research was enabled in part by support provided by Compute Ontario and the Digital Research Alliance of Canada.